SOME FUNCTION SPACES RELATED TO THE BROWNIAN MOTION ON SIMPLE NESTED FRACTALS

KATARZYNA PIETRUSKA-PALUBA

Abstract. In this paper we identify the domain of the Dirichlet form associated with the Brownian motion on simple nested fractals with an integral Lipschitz space. This result generalizes such an identification on the Sierpiński gasket, carried on by Jonsson.